2) 17pt) Solve the following IVP. Xi = -X1 + (3/2)x2 x1(2) = 1 xz = (-1/6)x1 - 2x2 X2(2) = 0
Answer #1

*=-x, + Zuz - 1/2 I x - 28 X= or 3/2/a 2)(*) X2 16 -2 let Az (132) 312 -1/6 of A equation The characterstic -1-1 1A-XI) = 3/10:18 =) 2, +322=0, Here Zz is huee let zz=1 =) Z,=-3 1-3 - be Case I ( g Iz for second A2-3 , let associated eigen rector 1/4 e 2t 0 kan ces t2cze 3 = 0 = £4, + 2 2 2 2 = -3 = y + 342=-6 let Y2=0 2) Y = -6 The general solution of system -3627 - 327